OpenStudy (anonymous):

2log6+log (1/3)=log12 which properties of logarithmics are/is used? power property and product property quotient property and product property quotient property only power property only

OpenStudy (anonymous):

Well If I asked you what 2log6 + log(1/3) was, how would you solve it (without using a calculator)?

OpenStudy (anonymous):

my teacher didnt go over this but still put it on the end of course exam

OpenStudy (anonymous):

log (1/3) is saying -log 3. With that we can use quotient property

OpenStudy (anonymous):

If it's just 2log6 +log(1/3) you can reduce it down to 2log6-log(3) because log(1) equal zero

OpenStudy (anonymous):

2log6 can be written as log6^2 so, log 36

OpenStudy (anonymous):

log36/log 3 is log 12

OpenStudy (anonymous):

No it's not. log36 - log3 is log12.

OpenStudy (anonymous):

same thing

OpenStudy (anonymous):

make sense?

OpenStudy (anonymous):

no it isn't. \[\frac{log\ 36}{log\ 3} \ne log(\frac{36}{3})\]

OpenStudy (anonymous):

log36-log3 =log36/log3

OpenStudy (anonymous):

i'm not saying log(36/3)

OpenStudy (anonymous):

No, it doesn't. log 36 - log 3 = log(36/3)

OpenStudy (anonymous):

>>> log(36, 10) - log(3, 10) 1.0791812460476247 >>> log(36,10)/log(3,10) 3.2618595071429146

OpenStudy (anonymous):

Very different numbers.

OpenStudy (anonymous):

try it in your calculator, (or wolframalpha)

OpenStudy (anonymous):

log(36/3) is log(12) i checked with calculator and got the same answer so sorry for confusion just disregard the log36/log3

OpenStudy (anonymous):

You can see quite trivially that it will be equal because 36/3 is 12. So log of 36/3 would be the log of 12. The problem is more about finding that log(36) - log(3) can be written as log(36/3)

OpenStudy (anonymous):

ok so what was the problem?

OpenStudy (anonymous):

The problem with your answer was that you said log(36) - log(3) = log(36)/log(3). But the simplest solution is to just say: \[2log6+log (1/3)\] \[=log36 + log(\frac{1}{3})\] \[=log (36 \times \frac{1}{3})\] \[=log(12)\] Which just uses the power and product rules. That said though you could also solve it by just using the product rule, or using the quotient rule and the power rule. So there's no non-ambiguous way to answer the multiple choice.